At a certain temperature, the equilibrium constant Kc for the reaction 2BrCl(g) ⇌ Br2(g) + Cl2(g) is 0.035. A mixture of 8.00 moles of Br2 and 12.00 moles of Cl2 are placed in a 20.0 L container is allowed to come to equilibrium. What is the equilibrium concentration of BrCl?
1
Expert's answer
2020-09-24T13:51:11-0400

2BrCl "\iff" Br2+Cl2

Kc= .035

Concentration of Br2= 8/20=.4

Concentration of Cl2=12/20=.6

2BrCl "\iff" Br2+ Cl2

Intial conc. [0] .4 .6

Change in. 2x -x -x

At equilibrium 2x .4-x .6-x

Conc.

Kc= [Br2] [Cl2] / [ BrCl2]

.035= (.4-x)(.6-x)/ (2x)2

solve for x

value of x is .682

Concentration of BrCl2= 2x=2×.682

= 1.364